# FENIX: Enabling In-Network DNN Inference with FPGA-Enhanced Programmable Switches

### v1

@@ Abstract
Machine learning (ML) is increasingly used in network data planes for advanced traffic analysis, but existing solutions (such as FlowLens, N3IC, BoS) still struggle to simultaneously achieve low latency, high throughput, and high accuracy. To address these challenges, we present FENIX, a hybrid in-network ML system that performs feature extraction on programmable switch ASICs and deep neural network inference on FPGAs. FENIX introduces a Data Engine that leverages a probabilistic token bucket algorithm to control the sending rate of feature streams, effectively addressing the throughput gap between programmable switch ASICs and FPGAs. In addition, FENIX designs a Model Engine to enable high-accuracy deep neural network inference in the network, overcoming the difficulty of deploying complex models on resource-constrained switch chips. We implement FENIX on a programmable switch platform that integrates a Tofino ASIC and a ZU19EG FPGA directly, and evaluate it on real-world network traffic datasets. Our results show that FENIX achieves microsecond-level inference latency and multi-terabit throughput with low hardware overhead, and delivers over 90% accuracy on mainstream network traffic classification tasks, outperforming SOTA.
@@ 1 Introduction
To address these latency challenges, researchers have explored moving ML inference closer to the data path. One promising approach is to implement inference directly on programmable devices such as SmartNICs, as demonstrated by systems like N3IC [40]. This offloads computation from endpoint CPUs and significantly reduces inference latency. However, new bottlenecks arise: for example, N3IC achieves a maximum throughput of only 40 Gbps, and even the latest commercial SmartNICs are limited to around 400 Gbps [32]. Such throughput is often insufficient for core networks, where switch ASICs handle 4 to 12 Tbps [22, 21], far exceeding current SmartNIC capabilities.
Recent advances in in-network ML have led researchers to implement ML models directly on programmable switch ASICs. For example, systems such as Leo [24] and NetBeacon [57] deploy decision tree and random forest models on Tofino switches using Match Action Tables (MAT), while BoS [51] demonstrates the feasibility of running binary recurrent neural networks (RNNs) on similar hardware. These approaches can potentially enable line-rate ML inference at significantly higher throughput. However, they are typically subject to stringent computational resource constraints, which limit their effectiveness for complex multi-classification tasks (see Section § 7 for details).
These observations raise several key questions: Why do existing switch ASICs struggle to support more sophisticated ML tasks? Beyond the programmable devices discussed above, what other platforms can effectively balance accuracy and latency for in-network ML? In this paper, we present FENIX (FPGA-Enabled Neural Inference eXecution for network switches), a system designed to achieve low latency, high throughput, and high accuracy for in-network intelligent traffic analysis.
Figure 1: Design sapce in Intelligent Network.
@@
FENIX consists of two main components: the Data Engine and the Model Engine, which together enable in-network traffic analysis on the data plane. The Data Engine extracts relevant features from high-throughput network traffic, while the Model Engine performs machine learning inference on these features. This modular design enables FENIX to achieve three key goals: low latency by avoiding software processing and enabling direct hardware interfacing, high throughput through efficient feature extraction and communication, and high accuracy by supporting full DNN models with minimal quantization loss.
Contributions. The main contribution of this paper is the design, implementation, and evaluation of FENIX, the first system to leverage FPGA augmentation to extend the capabilities of programmable switches for in-network machine learning. Over six months, we designed and manufactured a custom switch, integrating a programmable Tofino switch chip and a ZU19EG FPGA on Printed Circuit Board (PCB). We evaluate FENIX on challenging tasks, including VPN encrypted traffic classification and malware detection. Experimental results show that FENIX achieves up to 537$\times$ lower inference latency compared to control plane-based approaches, and improves classification accuracy by up to 21% over SOTA. Our prototype shows low hardware resource overhead, indicating that FPGA-enhanced switches are practical for deployment in high-speed network environments.
@@ 2 BACKGROUND AND MOTIVATION
Motivation. Current research in-network traffic analysis mainly follows two directions: adapting machine learning models to fit data plane constraints (e.g., transforming decision trees into match-action table representations) [51, 57, 24, 49, 26], or sampling data for offline analysis on servers [10, 55, 25]. However, both approaches inevitably compromise either model accuracy or introduce significant latency, and thus fall short of modern network requirements. The core challenge is to deploy powerful deep neural networks (DNNs) in high-speed network environments. While DNNs excel at traffic classification and anomaly detection, they face a critical bottleneck: programmable switches offer limited computational resources and restricted instruction sets, making complex DNN execution infeasible; meanwhile, hardware accelerators such as FPGAs, though efficient for neural network computation, provide processing bandwidth orders of magnitude lower than the Tbps-level throughput of switches.
@@ 3 FENIX Design Overview
Figure 2 shows the architecture of FENIX, which is designed to address two main challenges in integrating deep learning inference with programmable switches. The first challenge is managing the throughput mismatch between programmable switching ASICs and FPGAs under dynamic and bursty network traffic. The second challenge is enabling efficient deployment of deep neural network inference within the constraints of switch-based environments.
To address the throughput mismatch, FENIX uses a Data Engine that manages feature caching and transmission control. The Data Engine consists of a Flow Tracker, Rate Limiter, and Buffer Manager, which together dynamically determine when flows’ features are transmitted to the FPGA. By coordinating these components, the Data Engine adapts to changing traffic conditions and ensures balanced utilization of both the switch ASIC and FPGA.
To support DNN inference on the data path, FENIX employs a Model Engine responsible for both flow identification and model execution. The Model Engine integrates a Vector I/O Processor with a DNN Inference Module, and supports a variety of model structures, including embedding, fully connected, convolutional, and recurrent layers. By synthesizing model parameters and feature dimensions at design time, FENIX efficiently utilizes FPGA resources while maintaining flexibility for diverse neural network architectures.
With this design, FENIX achieves high-throughput, low-latency, and accurate DNN inference for in-network traffic analysis. The following sections detail the design of each module.
@@ 4 Data Engine
As shown in Figure 2, the Data Engine of FENIX is implemented on programmable switch ASICs and consists of three key modules working in concert. The Flow Tracker maintains complex hash table structures to accurately record state information for each network flow, including flow duration $T_{i}$ and packet count $C_{i}$, which serve as the basis for subsequent analysis. The Rate Limiter uses an enhanced token bucket algorithm to control the transmission frequency of feature information according to the computed probability function $P(T_{i},C_{i})$, effectively addressing throughput mismatch between the switch and FPGA, and dynamically adapting to the current traffic rate. This probabilistic sampling method, based on flow features, helps maintain system stability under high-traffic conditions while ensuring the quality of inputs for model inference. The Buffer Manager employs ring buffer techniques to temporarily store feature data, attaches the data to mirrored packet headers, and efficiently forwards traffic features to the Model Engine for further processing.
@@ 4.1 Flow Tracker
Flow Counting Mechanism. We define the start time of a flow as the instant when its first packet arrives at the switch, and count the number of flows that send packets within each timeout interval $T_{w}$. This statistical approach, which focuses on newly arrived flows within each interval, helps mitigate bias arising from missing the exact start or end times of flows [43]. The mechanism operates as illustrated in Figure 4 (a): the flow counter detects new flows by checking hash registers associated with each flow. Upon detecting a new flow, the counter increments $n$ by 1. At the end of each $T_{w}$ period, both the hash registers and the flow count $n$ are reset and subsequently recalculated by the control plane, ensuring accurate and up-to-date flow statistics.
Buffer Index Update. As depicted in Figure 4 (b), since the data plane cannot perform modulo operations directly, we maintain a separate buffer index (buff_idx) for each flow that increments with each packet and resets to 1 when reaching buffer size. The flow’s packet counter (Pkt_cnt) independently tracks total packets. This design enables a logical ring buffer where the buffer index cycles through available space, ensuring new data replaces old data in a circular manner.
@@ 4.2 Rate Limiter
The rate limiter maintains global traffic statistics within each timing window $T_{w}$, including Flow_cnt (the total number of flows) and Pkt_cnt (the total number of packets processed). Leveraging these statistics, the system achieves several advantages: high-speed flows are more likely to fail when requesting tokens, thereby preserving inference opportunities for lower-rate flows, and by capping the token bucket size to no more than the queue length, the system can effectively absorb traffic bursts without causing excessive queuing or packet drops.
@@
Based on these criteria, we propose the following piecewise probability model. The formal proof of its fairness can be found in Appendix A.
$P_{i}(T_{i},C_{i})=\begin{cases}\frac{C_{i}(VT_{i}-N)}{QT_{i}-NC_{i}},&\text{if }T_{i}\in\left[\frac{N}{V},\frac{QT_{i}}{C_{i}V}\right]\,\left(\frac{N}{V}<\frac{QT_{i}}{C_{i}V}\right)\\ \frac{T_{i}(VC_{i}-Q)}{NC_{i}-QT_{i}},&\text{if }T_{i}\in\left[\frac{QT_{i}}{C_{i}V},\frac{N}{V}\right]\,\left(\frac{N}{V}>\frac{QT_{i}}{C_{i}V}\right)\\ 1,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}\geq\frac{N}{V}\\ 0,&\text{if }QT_{i}=NC_{i}\text{ and }T_{i}<\frac{N}{V}\end{cases}$ (2)
Figure 6: Probability curves of token generation model.
@@
2:     $T_{last}\leftarrow T_{now}$, $gap\leftarrow 0$triangleright Initialize for first packet
@@
4:     $gap\leftarrow T_{now}-T_{last}$, $T_{last}\leftarrow T_{now}$triangleright Calculate time gap
@@
6:$rand\leftarrow$ Random(), $prob\leftarrow$ LookupProbability()
7:$bucket\leftarrow bucket+gap$triangleright Refill tokens
@@
10:         $bucket\leftarrow bucket-cost$triangleright Consume token
@@
Token Bucket Algorithm Implementation. In our data plane design, the rate limiter tracks three essential state variables within a custom timing window $T_{w}$: the number of active flows, the total packet count, and the last packet arrival time $T_{last}$. The details of our probabilistic token bucket mechanism are presented in Algorithm 1. Upon the arrival of each packet, the algorithm first computes the time interval between the current arrival time $T_{now}$ and $T_{last}$ to determine the appropriate token replenishment (lines 1–4). Next, a random number $rand$ is generated and compared against the probability value $prob$ retrieved from the precomputed lookup table (line 5). If the packet is probabilistically selected ($rand<prob$) and the token bucket contains enough tokens, the required token $cost$ is deducted from the bucket and the feature vector is transmitted (lines 6–8). Otherwise, if the selection fails or tokens are insufficient, the algorithm only updates the token count based on the elapsed $gap$ (lines 9–14).
@@ 4.3 Buffer Manager
[…]
@@ 5 Model Engine
The Model Engine is responsible for performing low-latency inference on traffic features extracted by the Data Engine and is implemented on an FPGA. As illustrated in Figure 8, the Model Engine comprises two main components: a Vector I/O Processor, which manages flow information, and a DNN Inference Module, which executes the model inference. Upon receiving packets from the Data Engine, the Vector I/O Processor separates each packet into a flow identifier and its corresponding feature vector. Flow identifiers are temporarily stored in the Flow Identifier Queue to preserve ordering, while feature vectors are forwarded to the DNN Inference Module for processing.
Once inference is complete, the results are placed in the Output Queue. Each inference result in the Output Queue is paired with the corresponding flow identifier retrieved from the front of the Flow Identifier Queue, forming a new packet with both fields. This packet is then transmitted back to the Data Engine, enabling the programmable switch to execute appropriate follow-up operations on the relevant flow based on the inference result. This approach maintains correct mapping between flows and their inference results throughout processing.
@@ 5.1 Vector I/O Processor
Figure 8: Workflow of Rate Limiter.
@@ 6 Implementation
Experimental Hardware. As shown in Figure 9, our experiments are conducted on a high-performance programmable switch platform that integrates both FPGA and Tofino chips on a single board. The hardware is implemented using a 22-layer high-performance printed circuit board (PCB), utilizing back-drilling technology to balance 100 Gbps high-speed interconnection performance and manufacturing costs. The switching system is equipped with a dedicated power control network, supporting a core operating current of up to 100 A and peak control currents up to 300 A, with precise impedance control to ensure reliable power delivery.
The overall design process included approximately half a month for requirement analysis, one month for schematic design, two months for PCB design, and one and a half months for PCB manufacturing, soldering, and prototype hardware debugging. The interface design and testing between FPGA and Tofino required an additional month, with software driver development and system-level integration carried out in parallel.
@@
    Control plane CPU: Programmed to manage the Tofino chip, responsible for configuration, monitoring, and real-time probability model adjustments.
@@
Model Training and Quantization. We trained several neural network architectures, including Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), for intelligent traffic analysis and classification tasks. Our approach leverages universal and protocol-agnostic traffic features: sequences of raw packet lengths and inter-packet arrival times, which capture the temporal patterns of network flows without requiring any protocol-specific information. For feature extraction, we implemented a software-based sliding window mechanism to process raw PCAP files and generate feature vectors, each consisting of a sequence of packet length and timing data from a flow, paired with the corresponding class label for supervised learning.
Our model architectures generally include normalization layers to standardize input features, followed by either convolutional layers (for CNNs) or recurrent units (for RNNs) to capture temporal dependencies. These are further connected to fully connected layers that map extracted features to output classes, with ReLU activation functions introducing non-linearity. To address class imbalance in-network traffic datasets, we applied both oversampling and undersampling techniques during preprocessing, thereby preventing overfitting to dominant classes and improving the representation of minority classes. This comprehensive approach results in more robust models with higher accuracy and better generalization for identifying diverse traffic types.
After training, we performed quantization using Vitis-AI [7] tools to convert floating-point model parameters to 8-bit integer (INT-8) format for efficient hardware deployment. The quantization process assigns different decimal point positions to different layers, based on their activation distributions, to preserve model accuracy. Our experimental evaluation shows that quantization substantially reduces computation and storage requirements while maintaining classification accuracy, with only negligible performance degradation. The quantized models are thus well-suited for deployment on resource-constrained FPGA platforms, enabling high-throughput, low-latency inference without sacrificing classification quality or robustness.

### v3

@@ Abstract
Machine learning (ML) is increasingly used in network data planes for advanced traffic analysis, but existing solutions (such as FlowLens, N3IC, BoS) still struggle to simultaneously achieve low latency, high throughput, and high accuracy. To address these challenges, we present FENIX, a hybrid in-network ML system that performs feature extraction on programmable switch ASICs and deep neural network inference on FPGAs. FENIX introduces a Data Engine that leverages a probabilistic token bucket algorithm to control the sending rate of feature streams, effectively addressing the throughput gap between programmable switch ASICs and FPGAs. In addition, FENIX designs a Model Engine to enable high-accuracy deep neural network inference in the network, overcoming the difficulty of deploying complex models on resource-constrained switch chips. We implement FENIX on a programmable switch platform that integrates a Tofino ASIC and a ZU19EG FPGA directly, and evaluate it on real-world network traffic datasets. Our results show that FENIX achieves microsecond-level inference latency and multi-terabit throughput with low hardware overhead, and delivers over 90% accuracy on mainstream network traffic classification tasks, outperforming the state of the art.
@@ 1 Introduction
To address these latency challenges, researchers have explored moving ML inference closer to the data path. One promising approach is implementing inference directly on programmable devices such as SmartNICs, as demonstrated by systems like N3IC [50]. This offloads computation from endpoint CPUs and significantly reduces inference latency. However, new bottlenecks arise: for example, N3IC achieves maximum throughput of only 40 Gbps, and even the latest commercial SmartNICs are limited to 400 Gbps [42]. Such throughput is often insufficient for core networks, where switch ASICs handle 4 to 12Tbps[31, 30], far exceeding current SmartNIC capabilities. While deploying multiple NICs can scale throughput, this approach introduces significant costs and new scheduling challenges, as packet scheduling consumes significant CPU resources [26].
Recent advances in in-network ML have led researchers to implement ML models directly on programmable switch ASICs. For example, systems such as Leo [33] and NetBeacon [71] deploy decision tree and random forest models on Tofino switches using Match Action Tables (MAT), while BoS [64] demonstrates the feasibility of running binary recurrent neural networks (RNNs) on similar hardware. These approaches can potentially enable line-rate ML inference at significantly higher throughput. However, they are typically subject to stringent computational resource constraints, which limit their effectiveness for complex multi-classification tasks (see § 7 for details).
These observations raise several key questions: Why do existing switch ASICs struggle to support more complex ML tasks? Beyond the programmable devices discussed above, what other platforms can balance accuracy and latency for in-network ML? In this paper, we present FENIX (FPGA-Enabled Neural Inference eXecution for network switches), a system designed to achieve low latency, high throughput, and high accuracy for in-network intelligent traffic analysis.
Figure 1: Design space in Intelligent Network.
@@
FENIX consists of two main components: the Data Engine and Model Engine, which enable in-network traffic analysis on the data plane. The Data Engine extracts features from high-throughput traffic, while the Model Engine performs ML inference on these features. This design enables FENIX to achieve three key goals: low latency by avoiding software processing, high throughput through efficient feature extraction, and high accuracy by supporting full DNN models with minimal quantization loss.
[…]
@@ 2 BACKGROUND AND MOTIVATION
Motivation. Current research in-network traffic analysis mainly follows two directions: adapting machine learning models to fit data plane constraints (e.g., transforming decision trees into match-action table representations) [64, 71, 33, 60, 35], or sampling data for offline analysis on servers [12, 69, 34]. However, both approaches inevitably compromise either model accuracy or introduce significant latency, and thus fall short of modern network requirements. The core challenge is to deploy powerful deep neural networks (DNNs) in high-speed network environments. While DNNs excel at traffic classification and anomaly detection, they face a critical bottleneck: programmable switches offer limited computational resources and restricted instruction sets, making complex DNN execution infeasible; meanwhile, hardware accelerators such as FPGAs, though efficient for neural network computation, provide processing bandwidth orders of magnitude lower than the Tbps-level throughput of switches. Even the most advanced FPGAs cannot directly process network traffic from switching chips. Therefore, it is necessary to select key feature information for the FPGA and implement rate control to adapt to different traffic rates.
@@ 3 FENIX Design Overview
Figure 2 shows the architecture of FENIX, which is designed to address two main challenges in integrating deep learning inference with programmable switches. The first challenge is enabling the data plane to execute deep neural network models for fine-grained traffic control. The second challenge is bridging the throughput gap between FPGAs and programmable switching ASICs.
To address the throughput mismatch, FENIX designs a Data Engine that manages feature caching and transmission control. The Flow Tracker is motivated by the need to enable per-flow precise inference, providing the foundation for applying DNN models to individual network flows. The Rate Limiter employs a token bucket mechanism with probabilistic control to regulate the rate at which each flow sends feature information, preventing the FPGA from being overwhelmed by excessive feature data. The Buffer Manager uses a ring buffer structure with mirrored headers, enabling the data engine to pass stateful feature information to the model engine for processing.
To support DNN inference on the data path, FENIX designs a Model Engine responsible for both flow identification and model execution. The Vector I/O Processor is designed to support high-throughput, low-latency data exchange between the data engine and model engine, and caches flow identifiers so that the ASIC can accurately identify inference results. The DNN Inference Module performs parallel deep neural network inference on feature vectors, maximizing FPGA resource utilization.
With this design, FENIX achieves high-throughput, low-latency, and accurate DNN inference for in-network traffic analysis. The following sections detail each module design.
@@ 4 Data Engine
As shown in Figure 2, the Data Engine of FENIX is implemented on programmable switch ASICs and consists of three key modules working in concert. The Flow Tracker maintains complex hash table structures to accurately record state information for each network flow, including flow duration $T_{i}$ and packet count $C_{i}$, which serve as the basis for subsequent analysis. The Rate Limiter uses an enhanced token bucket algorithm to control the transmission frequency of feature information according to the computed probability function $P(T_{i},C_{i})$, effectively addressing throughput mismatch between the switch and FPGA, and dynamically adapting to current traffic rate. This probabilistic sampling method, based on flow features, helps maintain system stability under high-traffic conditions while ensuring quality of inputs for model inference. To address FPGA inference latency that could potentially allow some malicious packets to pass undetected in per-packet inference scenarios, FENIX integrates lightweight decision trees as a complementary mechanism. For flows awaiting FPGA inference results, the system temporarily uses decision tree inference; once FPGA results become available, the flow processing strategy is immediately updated. The Buffer Manager employs ring buffer techniques to temporarily store feature data, attaches data to mirrored packet headers, and efficiently forwards traffic features to the Model Engine for further processing.
@@ 4.1 Flow Tracker
Flow Counting Mechanism. We define the start time of a flow as the instant when its first packet arrives at the switch, and count the number of flows that send packets within each timeout interval $T_{w}$. This statistical approach, which focuses on newly arrived flows within each interval, helps mitigate bias arising from missing the exact start or end times of flows [53]. The mechanism operates as illustrated in Figure 4 (1): the flow counter detects new flows by checking hash registers associated with each flow. Upon detecting a new flow, the counter increments $n$ by 1. At the end of each $T_{w}$ period, both the hash registers and the flow count $n$ are reset by the control plane to begin a new counting cycle, ensuring accurate and up-to-date flow statistics.
Buffer Index Update. As depicted in Figure 4 (2), since the data plane cannot perform modulo operations directly, we maintain a separate buffer index (buff_idx) for each flow that increments with each packet and resets to 1 when reaching buffer size. The flow’s packet counter (Pkt_cnt) independently tracks total packets. This design enables a logical ring buffer where the buffer index cycles through available space, ensuring new data replaces old data in a circular manner.
@@ 4.2 Rate Limiter
The rate limiter maintains global traffic statistics within each timing window $T_{w}$, including Flow_cnt (the total number of flows) and Pkt_cnt (the total number of packets processed). While standard work-conserving algorithms such as WFQ (Weighted Fair Queuing) are mature solutions for ensuring inter-flow fairness, our goal is to achieve precise rate control based on per-flow state. In practice, it is infeasible to allocate physical queues for each flow, and alternative approximation schemes cannot guarantee precise rate matching between FPGA and programmable switches. Our probabilistic token bucket mechanism provides a lightweight, hardware-friendly alternative that can be efficiently implemented on existing switch architectures. Leveraging these statistics, the system achieves several advantages: high-speed flows are more likely to fail when requesting tokens, thereby preserving inference opportunities for lower-rate flows, and by capping the token bucket size to no more than the queue length, the system can effectively absorb traffic bursts without causing excessive queuing or packet drops.
Figure 6: Probability curves of token generation model.
@@
Based on these criteria, we propose the following piecewise probability model.
[…]
@@
2:  $T_{last}\leftarrow T_{now}$, $gap\leftarrow 0$ $\triangleright$ Initialize for first packet
@@
4:  $gap\leftarrow T_{now}-T_{last}$, $T_{last}\leftarrow T_{now}$ $\triangleright$ Calculate time gap
@@
[…]
@@
10:   $bucket\leftarrow bucket-cost$ $\triangleright$ Consume token
@@
Token Bucket Algorithm Implementation. In our data plane design, the rate limiter tracks three essential state variables within a custom timing window $T_{w}$: the number of active flows, the total packet count, and the last packet arrival time $T_{last}$. The details of our probabilistic token bucket mechanism are presented in Algorithm 1. Upon the arrival of each packet, the algorithm first computes the time interval between the current arrival time $T_{now}$ and $T_{last}$ to determine the appropriate token replenishment (lines 1–4). Next, a random number randrand is generated and compared against the probability value probprob retrieved from the precomputed lookup table (line 5). If the packet is probabilistically selected ($rand<prob$) and the token bucket contains enough tokens, the required token costcost is deducted from the bucket and the feature vector is transmitted (lines 6–8). Otherwise, if the selection fails or tokens are insufficient, the algorithm only updates the token count based on the elapsed gapgap (lines 9–14).
@@ 4.3 Buffer Manager
Ring Buffer Design. Inspired by BoS [64], the Buffer Manager assigns a dedicated ring buffer to each network flow, which temporarily stores feature vectors until they are processed by the Model Engine. For each flow, the ring buffer holds features extracted from preceding packets ($F_{1}$$F_{8}$), including packet intervals and packet lengths, as well as the current packet’s feature ($F_{9}$, stored in metadata). This architecture allows the system to preserve the temporal and sequential relationships between packets within the same flow, offering a comprehensive context for subsequent analysis. When a buffer reaches capacity, it operates in a circular (FIFO) manner, overwriting the oldest features with new entries. This ensures that the buffer consistently reflects the most recent traffic characteristics, efficiently utilizes memory, and maintains the necessary flow context for downstream processing.
@@ 5 Model Engine
The Model Engine performs low-latency inference on traffic features from the Data Engine and is implemented on an FPGA. As shown in Figure 8, it comprises two components: a Vector I/O Processor which manages flow-info and a DNN Inference Module which executes model inference. Upon receiving packets, the Vector I/O Processor separates each into flow identifier and feature vector. Flow identifiers are stored in the Flow Identifier Queue to preserve ordering, while feature vectors are forwarded to DNN Inference Module.
After inference completion, the results are placed in the Output Queue. Each inference result in the Output Queue is paired with the corresponding flow identifier retrieved from the front of the Flow Identifier Queue, forming a new packet with both fields. This packet is then transmitted back to the Data Engine, enabling the switch chip to execute follow-up operations on the relevant flow based on the inference result. This approach maintains correct mapping between flows and their inference results throughout processing.
@@ 5.1 Vector I/O Processor
Figure 8: Workflow of Model Engine.
@@ 6 Implementation
Experimental Hardware. As shown in Figure 9, our experiments are conducted on a high-performance programmable switch platform that integrates both FPGA and Tofino chips on a single board. The hardware is implemented using a 22-layer high-performance printed circuit board (PCB), utilizing back-drilling technology to balance 100 Gbps high-speed interconnection performance and manufacturing costs. The switching system is equipped with a dedicated power control network, supporting a core operating current of up to 100 A and peak control currents up to 300 A, with precise impedance control to ensure reliable power delivery. The overall design process included approximately half a month for requirement analysis, one month for schematic design, two months for PCB design, and one and a half months for PCB manufacturing, soldering, and prototype hardware debugging. The interface design and testing between FPGA and Tofino required an additional month, with software driver development and system-level integration carried out in parallel.
@@
    Control plane CPU: Programmed to manage the Tofino chip, responsible for configuration, monitoring, and probability model adjustments.
@@
Model Training and Quantization. We trained CNNs and RNNs for traffic classification using protocol-agnostic features: packet lengths and inter-packet arrival times. These features capture temporal patterns without requiring protocol-specific information. Our models include normalization layers, convolutional/recurrent units for temporal dependencies, and fully connected layers with ReLU activation. To address class imbalance, we applied oversampling and undersampling techniques during preprocessing.
After training, we used Vitis-AI [7] to quantize models from floating-point to INT-8 format for efficient FPGA deployment. The quantization process assigns different decimal positions to layers based on activation distributions to preserve accuracy. Our evaluation shows that quantization substantially reduces computation and storage requirements while maintaining classification performance with negligible degradation, enabling high-throughput, low-latency inference on resource-constrained FPGA platforms.